OpenStudy (alexh107):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1^2 + 4^2 + 7^2 + ... + (3n-2)^2 = n(6n^2 - 3n -1)/2

OpenStudy (alexh107):

For n = 1 I got \[1^2 = 1\] after simplifying but do not know where to go from there.

OpenStudy (welshfella):

next assume that the statement is true for n = k and write the formula for n = k Then find a formula for the (k + 1)th term in the series and add it to the statement for n = k Then try to manipulate yhe latter to see if you can get it in the same form as for n = k with the k replaced by k+1. If you can do this then you have proved the statement

OpenStudy (alexh107):

I'm confused as to what the k is representing. I know to replace n with k but I don't understand what that mean.s

OpenStudy (welshfella):

sum for n = k is k(6k^2 - 3k -1)/2 k represents any integer value of n

OpenStudy (welshfella):

(k + 1)th term = 3(k + 1) - 2)^2

OpenStudy (alexh107):

Oh okay, so for n = k to prove it is all I have to do is replace n with k? Are there any other steps? And with K+1 is it the same thing, just replacing n with k+1?

OpenStudy (welshfella):

You add the (k+1) term to the formula for k:- Sum of k+1 terms = n(6n^2 - 3n -1)/2 + (3k - 1)^2

OpenStudy (welshfella):

Sorry that should be k(6k^2 - 3k -1)/2 + (3k - 1)^2

OpenStudy (alexh107):

Oh okay, that makes sense. Are there any other steps to prove it or is it already proved?

OpenStudy (welshfella):

no This is the hard part. You need to convert this to 'the formula for k' replaced by k+1 in other words something like (k + 1)(6(k + 1)^2 - 3(k+1) -1) / 2 - do you see the pattern - the 'k' in original formula is rplaced by k+1

OpenStudy (alexh107):

What I have so far: For n = 1 \[1^2 = 1(6 \times 1^2 - 3 \times 1 - 1)/2 = 1\] For n = k \[k(6k^2 -3k - 1)/2\] For k+1 \[(3(k+1)-2)^2 + k(6k^2-3k-1)/2\]

OpenStudy (welshfella):

yes ths correct so far

OpenStudy (alexh107):

Okay I see that the k was replaced by k+1, I don't know what to do next though.

OpenStudy (welshfella):

no its a tricky on because the trinomial won;t factor I'll have a look at it..

OpenStudy (alexh107):

In the directions it says the statement can be false, I don't know if that helps.

OpenStudy (welshfella):

Well thats what i think . I cant see how we can manipulate that to get the required expression.

OpenStudy (alexh107):

The only problem is I have only two questions and the other one is already false so I'm doubtful they would give me two false questions.

OpenStudy (welshfella):

Lets ee if its tru for n = 3 3(6*9 - 9 - 1) / 2 = 66 yes its true for n = 3 thats not a proof though

OpenStudy (alexh107):

So it is true we just need to figure out how to simplify it to get it the n=k statement?

OpenStudy (welshfella):

if its true we will get the sum of k statement with the k replaced by k+1. But the algebra is a bit hairy!

OpenStudy (alexh107):

Okay, I think I'll work on it for a bit. Thank you for all your help.

OpenStudy (welshfella):

I guess you could simplify your expression to a cubic polynomial. Then you do the same with the original expressiion with the n replaced by k + 1. If these 2 agree then it is proved.

OpenStudy (welshfella):

yes - I got a result that way But i cant tell you if its true or false.